OpenStudy (anonymous):

jack is an actuary working for a corporate pension fund. He needs to have $14.6 million grow to $22 million in six years. What interest rate compounded annually does he need for this investment?

OpenStudy (amistre64):

just need to solve the formula ...

OpenStudy (anonymous):

don't know what to do now

OpenStudy (amistre64):

what is our compound interest formula ... oh good, dint even need to ask

OpenStudy (amistre64):

A = P(1+r)^t A/P = (1+r)^t log(A/P) = t log(1+r) log(A/P) ------- = t log(1+r)

OpenStudy (amistre64):

pfft, looking to solve for r, not t

OpenStudy (amistre64):

root it

OpenStudy (amistre64):

A/P = (1+r)^t (A/P)^(1/t) = 1+r (A/P)^(1/t) - 1= r

OpenStudy (anonymous):

got it

OpenStudy (amistre64):

halfway thru i forgot what we were solving for :)